OpenStudy (anonymous):

Find the slope and y intercepts of the line... 4x-3y=36?

OpenStudy (anonymous):

substitute 0 for x to find y-intercept 4(0)-3y=36 -3y=36 y=-12 substitute 0 for y to find x-intercept 4x-3(0)=3y 4x=36 x=9 so the x- and y-intercepts are (9,0) and (0,-12)